# EDNet: A Distortion-Agnostic Speech Enhancement Framework with Gating Mamba Mechanism and Phase Shift-Invariant Training

### v1

@@ V-C GM Module Ablation Study
TABLE V: Results of GM Module Ablation Study.
@@
To validate the design choices of the GM module, we perform an ablation study by systematically disabling or modifying its internal components. The results are summarized in Table V. We first evaluate the impact of removing either the masking or mapping branch. Removing the masking branch results in moderate performance degradation (row A), while eliminating the mapping branch leads to a substantial drop across all metrics (row B). This highlights the importance of the mapping branch in generative tasks, such as bandwidth extension, where reconstructing missing spectral content is essential. In contrast, the masking branch helps preserve fine-grained details and performs well in scenarios where most of the original signal remains intact but is locally degraded, such as in denoising and dereverberation. However, under more challenging conditions involving simultaneous noise, reverberation, and bandwidth limitation, masking alone proves inadequate, emphasizing the necessity of a generative component. Next, we evaluate whether a simple combination of the two branches without dynamic gating can effectively leverage their strengths (row C). When masking and mapping outputs are averaged using a fixed fusion weight of 0.5, performance improves over using masking alone but falls short of the mapping-only baseline. This suggests that static fusion mechanism can result in an averaging effect, rather than a meaningful integration of complementary features.
@@
Taken together, these results show that the GM module is not merely an ensemble of two strategies, but a tightly integrated design. Optimal performance depends critically on how gating is performed, where it is positioned, and how the branches interact through shared and conditional information.
@@ V-D1 Task-Specific Behavior
TABLE VI: Mean Gate Values for Task-specific Scenarios.
@@
TABLE VII: Mean Gate Values for Input-specific Scenarios.
@@
Figure 4: Visualization of gate maps in input-specific scenarios using the multi-distortion model. Gate maps of the multi-distortion model given input of (a) additive noise, (b) reverberation, (c) 8 kHz bandwidth limitation, and (d) all three distortions combined.
Table VI summarizes the average gate activations for models trained on individual distortion types. The denoising model exhibits the strongest masking tendency (0.772), followed by the bandwidth extension models at 8 kHz (0.760) and 4 kHz (0.731). The dereverberation model shows the lowest masking ratio (0.693), reflecting increased reliance on mapping. These trends align with the nature of each task. In denoising, where speech remains largely intact and noise is additive, masking effectively suppresses interference. Dereverberation introduces globally diffuse and smeared distortions, necessitating broader reconstruction through mapping. Bandwidth extension displays more nuanced behavior: as input bandwidth narrows, the model increasingly depends on mapping to infer missing high-frequency components, while preserving low-frequency content through masking.
Fig. 3 visualizes this behavior. Under additive noise and reverberation, the gating pattern remains broadly uniform; however, reverberation yields slightly lower gate values, reflecting a modestly greater reliance on mapping compared to denoising. For bandwidth-limited inputs, the gate map reveals strong frequency dependency: masking dominates in the lower band and gradually diminishes at higher frequencies, where reconstruction is required. This pattern resembles prior approaches that explicitly reuse the low-frequency band [23] or apply residual learning [21, 22]; however, in our case, the division emerges organically without any frequency-specific architectural design. The model learns to exploit the reliability of the lower band and invoke mapping only where necessary, suggesting that the gating mechanism adaptively modulates its behavior in response to match task requirements.
@@ V-D2 Input-Specific Behavior
We next assess whether the gating mechanism in the multi-distortion model remains adaptive to input characteristics despite being jointly trained. Table VII reports average gate values for test subsets grouped by distortion type. As expected, masking weights decrease with reduced bandwidth, from 0.594 at 8 kHz to 0.569 at 2 kHz, indicating an adaptive shift toward mapping. In contrast, variations in SNR have little effect on gating behavior, showing that masking remains robust across different noise intensities.
Figure 4 shows gate maps for both individual and combined distortions. Compared to single-task models, the multi-distortion model produces more complex spatial patterns, likely reflecting its exposure to overlapping distortion types during training. Despite this complexity, distinct gating behaviors remain evident: additive noise and reverberation lead to broadly distributed masking, while bandwidth limitation induces stronger mapping activity in high-frequency regions. Notably, under bandwidth-only conditions—where the low-frequency band remains clean—the model generates fine-grained masking peaks that align with harmonic structures. This behavior, not seen in single-task models, suggests that the gating module not only differentiates distortion types but also adapts to perceptually important signal characteristics.
@@ V-D3 Summary and Implications
V-E Phase Analysis and Impact of PSIT
@@
V-E1 Behavior of Phase Loss under Phase Shifts
@@
To examine how models receive supervision under phase misalignment, we conduct a controlled experiment using clean utterances from the VoiceBank+DEMAND dataset [60]. Synthetic phase shifts are applied to the phase spectrogram using Equation (1), followed by re-wrapping the phase values into the range $[-\pi,\pi]$ to maintain valid representation. We then compute the loss against the original (unshifted) reference using standard phase loss [10], both with and without PSIT, across various search grid configurations.
As shown in Fig. 5, the standard phase loss without PSIT is minimized only when the prediction is perfectly aligned, and it increases sharply with even slight misalignments, exhibiting fluctuations due to phase wrapping. In contrast, PSIT maintains near-zero loss over a wider range, determined by the selected search grid. For instance, a grid of $\{-0.5,0,0.5\}$ yields a zero-loss region spanning $[-1,1]$, based on the validity of linear approximation within ±$\pm 0.5$ of each grid point. More importantly, beyond this range, PSIT produces a less fluctuating and more consistent loss increase, avoiding the abrupt spikes observed with the standard phase loss. These behavior suggests that PSIT may offer a more stable and interpretable learning signal, even when the predicted phase is shifted yet structurally correct.
V-E2 Perceptual Metric Sensitivity to Phase Shifts
TABLE VIII: Impact of phase shift on objective speech quality metrics.
@@ V-E2 Perceptual Metric Sensitivity to Phase Shifts
$n=0.33$ 4.644 5.000 4.977 5.000 1.000
$n=0.66$ 4.643 5.000 4.769 5.000 1.000
$n=1$ 4.643 5.000 4.627 5.000 1.000
$n=1.33$ 4.642 5.000 4.541 5.000 1.000
$n=1.66$ 4.641 5.000 4.466 5.000 1.000
$n=1$1 frame 4.644 5.000 4.868 5.000 1.000
$n=1.33$1 frame 4.643 5.000 4.686 5.000 1.000
$n=1.66$1 frame 4.642 5.000 4.583 5.000 1.000
To assess how phase shifts influence standard perceptual objective metrics, we conduct an analysis using the clean set from the VoiceBank+DEMAND dataset [60]. Phase shifts of varying degrees are applied to the clean phase spectrogram in the STFT domain, followed by inverse STFT to reconstruct the waveform. The resulting signals are then evaluated using PESQ, CSIG, COVL, CBAK, and STOI [66], as summarized in Table VIII. We observe that PESQ exhibits only slight degradation with increasing phase shift, while CSIG, COVL, and STOI remain virtually unchanged across all shift levels. The only metric that consistently declines is CBAK, which is known to be sensitive to sample-level misalignment and likely interprets the resulting waveform discrepancies caused by a temporal shift as background noise artifacts.
@@
V-E3 Phase Reconstruction with and without PSIT
@@
For assessing the impact of PSIT on phase reconstruction quality, we conduct a controlled experiment where the model is trained to predict phase from clean ground-truth magnitude. This setup isolates the effect of phase prediction by removing confounding factors related to magnitude estimation. We compare the standard phase loss [10] (w/o PSIT) against four PSIT variants with increasing search grid sizes, each trained 15 times with different random seeds to ensure statistical reliability.
@@
V-E4 Training Dynamics in Multi-distortion Setting
@@
V-E5 Summary and Implications
In summary, our analyses confirm that PSIT effectively fulfills its goal of making phase supervision more tolerant to perceptually irrelevant shifts, both conceptually and empirically. It improves phase reconstruction quality, reduces penalties from benign misalignments, and facilitates more efficient optimization. As a rule-based strategy, PSIT introduces only negligible computational overhead during training. It thus offers a simple and lightweight addition to the training process that yields measurable gains in both performance and training efficiency, without requiring any architectural modifications.
@@ VII Conclusion
This paper proposes EDNet, a distortion-agnostic speech enhancement framework designed to address diverse distortion types through two key components: a gated integration of masking and mapping operations, and a phase-shift-invariant training (PSIT) objective. The GM module enables the adaptive fusion of complementary enhancement strategies, while PSIT introduces robustness to temporal misalignment in phase supervision. Together, these components form a flexible and generalizable architecture for speech enhancement. We validate the proposed design through comprehensive evaluations on single- and multi-distortion tasks, where EDNet achieved state-of-the-art or comparable performance without task-specific architectural adaptation. Ablation studies and gating behavior analyses confirm that the GM module not only improves performance but also allows the model to learn task- and input-specific control over its enhancement strategy. In addition, PSIT improves phase reconstruction quality, facilitates learning dynamics, and improve overall model performance by softening the effects of strict alignment constraints without an overhaed during training. These results highlight that integrating structural adaptability with shift-tolerant supervision offers a promising path toward building distortion-agnostic speech enhancement systems suited for real-world deployment

### v2

@@ V-C GM Module Ablation Study
TABLE VI: Results of GM Module Ablation Study.
@@
To validate the design choices of the GM module, we perform an ablation study by systematically disabling or modifying its internal components. The results are summarized in Table VI. We first evaluate the impact of removing either the masking or mapping branch. Removing the masking branch results in moderate performance degradation (row A), while eliminating the mapping branch leads to a substantial drop across all metrics (row B). This highlights the importance of the mapping branch in generative tasks, such as bandwidth extension, where reconstructing missing spectral content is essential. In contrast, the masking branch helps preserve fine-grained details and performs well in scenarios where most of the original signal remains intact but is locally degraded, such as in denoising and dereverberation. However, under more challenging conditions involving simultaneous noise, reverberation, and bandwidth limitation, masking alone proves inadequate, emphasizing the necessity of a generative component. Next, we evaluate whether a simple combination of the two branches without dynamic gating can effectively leverage their strengths (row C). When masking and mapping outputs are averaged using a fixed fusion weight of 0.5, performance improves over using masking alone but falls short of the mapping-only baseline. This suggests that heuristic fusion mechanism can result in an averaging effect, rather than a meaningful integration of complementary features.
@@
Taken together, these results show that the GM module is not merely an ensemble of two strategies, but a tightly integrated design. Optimal performance depends critically on how fusion is performed, where it is positioned, and how the branches interact through shared and conditional information.
@@ V-D1 Task-Specific Behavior
TABLE VII: Mean Gate Values for Task-specific Scenarios.
@@
TABLE VIII: Mean Gate Values for Input-specific Scenarios.
@@
Figure 4: Visualization of gate maps in input-specific scenarios using the multi-distortion model. Gate maps and output magnitude spectrogram of the multi-distortion model given input of (a) additive noise, (b) reverberation, (c) 8 kHz bandwidth limitation, and (d) all three distortions combined.
Table VII summarizes the average gate activations for models trained on individual distortion types. The denoising model exhibits the strongest masking tendency (0.772), followed by the bandwidth extension models at 8 kHz (0.760) and 4 kHz (0.731). The dereverberation model shows the lowest masking ratio (0.693), reflecting increased reliance on mapping. These trends align with the nature of each task. In denoising, where speech remains largely intact and noise is additive, masking effectively suppresses interference. Dereverberation introduces globally diffuse and smeared distortions, necessitating broader reconstruction through mapping. Bandwidth extension displays more nuanced behavior: as input bandwidth narrows, the model increasingly depends on mapping to infer missing high-frequency components, while preserving low-frequency content through masking.
Fig. 3 visualizes this behavior. Under additive noise and reverberation, the gating pattern remains broadly uniform; however, reverberation yields slightly lower gate values, reflecting a modestly greater reliance on mapping compared to denoising. For bandwidth-limited inputs, the gate map reveals strong frequency dependency: masking dominates in the lower band and gradually diminishes at higher frequencies, where reconstruction is required. This pattern resembles prior approaches that explicitly reuse the low-frequency band [liu2022neural] or apply residual learning [wang2021towards, lu2024towards]; however, in our case, the division emerges organically without any frequency-specific architectural design. The model learns to exploit the reliability of the lower band and invoke mapping only where necessary, suggesting that the gating mechanism adaptively modulates its behavior in response to match task requirements.
@@ V-D2 Input-Specific Behavior
We next assess whether the gating mechanism in the multi-distortion model remains adaptive to input characteristics despite being jointly trained. Table VIII reports average gate values for test subsets grouped by distortion type. As expected, masking weights decrease with reduced bandwidth, from 0.594 at 8 kHz to 0.569 at 2 kHz, indicating an adaptive shift toward mapping. In contrast, variations in SNR have little effect on gating behavior, showing that masking remains robust across different noise intensities.
Fig. 4 presents the gate maps and enhanced magnitude spectrograms for both individual and combined distortions. Compared to single-task models, the multi-distortion model produces more complex spatial patterns, likely reflecting its exposure to overlapping distortion types during training. Despite this complexity, distinct gating behaviors remain evident: additive noise and reverberation lead to broadly distributed masking, while bandwidth limitation induces stronger mapping activity in high-frequency regions. Notably, under single bandwidth limitation condition (Fig. 4c)—a scenario to which the multi-distortion model was not explicitly exposed during training—the module still generates fine-grained masking peaks that align with harmonic structures. The output spectrograms verify that these gating decisions facilitate effective signal recovery even in such unseen conditions; when the low-frequency band remains clean, the model reconstructs a high-frequency energy distribution that closely approximates the ground-truth reference. In multi-distortion cases (Fig. 4d), however, the model invokes a more intricate combination of masking and mapping to restore corrupted low-frequency regions. While this yields favorable results, the inherent difficulty of low-frequency reconstruction leads to subtle deviations in the high-frequency energy, potentially acting as a primary source of high-frequency artifacts. Overall, these results confirm that the gating module not only differentiates distortion types but also adaptively modulates its strategy based on the spectral characteristics of the input signal.
@@ V-D3 Summary and Implications
TABLE IX: Comparison of Parameter counts and FLOPs.
Method Param. Encoder Core Module Decoder Total
CMGAN[abdulatif2024cmgan] 1.83M 16.40G 29.20G 17.70G 63.30G
SEMamba [chao2024investigation] 2.25M 24.36G 16.00G 24.60G 64.96G
MP-SENet [lu2023explicit] 2.26M 16.40G 52.40G 17.72G 86.52G
EDNet 2.67M 20.52G 31.04G 11.10G 62.66G
- Mag stream 2.28M 16.40G 27.08G 8.86G 52.34G
- Pha stream 0.39M 4.12G 3.96G 2.24G 10.32G
V-E Model Complexity Analysis
In terms of computational cost, EDNet differs from prior DenseNet-based encoder–decoder models in two respects: (i) a dual-stream architecture that processes magnitude and phase separately, and (ii) stacked Gating Mamba (GM) blocks, each comprising two processing modules—masking and mapping. Managing the associated overhead is a primary concern in our model design. In this section, we analyze the design choices intended to address this overhead and their effects, from theoretical and empirical perspectives.
Table IX summarizes the parameter counts and FLOPs of EDNet and comparable DenseNet-based systems. While EDNet introduces a modest increase in parameters, it attains the lowest total FLOPs. These efficiency gains arise from three design choices. First, to offset the doubled computation incurred by executing both masking and mapping process, we select the TF-Mamba [chao2024investigation] module for its superior computational efficiency relative to Transformer-based alternatives. Second, whereas most DenseNet-based baselines employ four processing blocks, EDNet uses three GM blocks. Third, to minimize the overhead introduced by the dual-stream design, we allocate the majority of the capacity to the magnitude stream and make the phase stream deliberately lightweight by reducing its channel width to 32 (0.39M parameters; 10.32G FLOPs). This magnitude-dominant allocation is motivated by our phase-reconstruction analysis in following Sec. V-F2. We find that providing clean magnitude features enables the model to recover phase near ground-truth quality, indicating that accurate magnitude restoration is more critical to overall enhancement performance.
(a) Real-time Factor
(b) Memory Usage
Figure 5: Empirical efficiency comparison among CMGAN, SEMamba, MP-SENet and EDNet. (a) Real-time factor (RTF) and (b) GPU memory usage with increasing input duration.
In addition to the theoretical complexity analysis, we assess empirical efficiency of EDNet under real-world inference conditions. We measure the average real-time factor (RTF) and GPU memory consumption using 100 audio samples with durations from 1 to 30 seconds in 2-second increments. All experiments are conducted on an AMD EPYC 7543 CPU and a single NVIDIA A6000 GPU (48 GB VRAM). As shown in Fig. 5, SEMamba [chao2024investigation] attains the lowest RTF and memory usage, with EDNet following a similar trend. Despite having the lowest FLOPs, EDNet shows a measured RTF that is approximately 0.01 higher and consumes slightly more memory, likely due to the heavier core module and serial magnitude-to-phase reconstruction process.
When compared with Transformer-based counterparts such as CMGAN [abdulatif2024cmgan] and MP-SENet [lu2023explicit], the Mamba-based models (EDNet and SEMamba) show clear advantages on longer sequences, maintaining relatively stable RTF and memory usage across input durations. CMGAN and MP-SENet begin to lag behind EDNet at approximately 9 and 15 seconds, respectively, with both RTF and memory consumption rising sharply as input length increases. While the Mamba-based models keep memory usage under 16 GB even for 30-second audio, their processing limits on the 48 GB A6000 are reached at around 19 and 27 seconds, respectively, underscoring the scalability gap between Transformer- and Mamba-based architectures. These findings are consistent with prior studies showing that Mamba architectures scale more efficiently than Transformer-based designs for long-context processing [gu2024mambalineartimesequencemodeling, jiang2025speech].
In summary, key design choices of EDNet enable higher enhancement performance and broader applicability, but they can also increase computational demand. By adopting TF-Mamba, reducing the number of GM blocks, and allocating capacity toward the magnitude stream, we control the resulting overhead so that both theoretical and measured complexities remain comparable to DenseNet-based baselines. Across all tested input lengths, EDNet maintains a real-time factor below 0.04 with practical memory usage, meeting real-time requirements and scaling to long contexts.
V-F Phase Analysis and Impact of PSIT
@@
V-F1 Behavior of Phase Loss under Phase Shifts
@@
To examine how models receive supervision under phase misalignment, we conduct a controlled experiment using clean utterances from the VoiceBank+DEMAND dataset [valentini2016investigating]. Synthetic phase shifts are applied to the phase spectrogram using Equation (1), followed by re-wrapping the phase values into the range $\pi,\pi]$ to maintain valid representation. We then compute the loss against the original (unshifted) reference using standard phase loss [lu2023explicit], both with and without PSIT, across various search grid configurations.
As shown in Fig. 6, the standard phase loss without PSIT is minimized only when the prediction is perfectly aligned, and it increases sharply with even slight misalignments, exhibiting fluctuations due to phase wrapping. In contrast, PSIT maintains near-zero loss over a wider range, determined by the selected search grid. For instance, a grid of $\{-0.5,0,0.5\}$ yields a zero-loss region spanning [1,1][-1,1], based on the validity of linear approximation within $\pm 0.5$ of each grid point. More importantly, beyond this range, PSIT produces a less fluctuating and more consistent loss increase, avoiding the abrupt spikes observed with the standard phase loss. These behavior suggests that PSIT may offer a more stable and interpretable learning signal, even when the predicted phase is shifted yet structurally correct.
V-F2 Perceptual Metric Sensitivity to Phase Shifts
TABLE X: Impact of phase shift on objective speech quality metrics.
@@ V-F2 Perceptual Metric Sensitivity to Phase Shifts
$n=0.33$ - 4.644 5.000 4.977 5.000 1.000
$n=0.66$ - 4.643 5.000 4.769 5.000 1.000
$n=1$ - 4.643 5.000 4.627 5.000 1.000
$n=1.33$ - 4.642 5.000 4.541 5.000 1.000
$n=1.66$ - 4.641 5.000 4.466 5.000 1.000
$n=1$ -1 frame 4.644 5.000 4.868 5.000 1.000
$n=1.33$ -1 frame 4.643 5.000 4.686 5.000 1.000
$n=1.66$ -1 frame 4.642 5.000 4.583 5.000 1.000
To assess how phase shifts influence standard perceptual objective metrics, we conduct an analysis using the clean set from the VoiceBank+DEMAND dataset [valentini2016investigating]. Phase shifts of varying degrees are applied to the clean phase spectrogram in the STFT domain, followed by inverse STFT to reconstruct the waveform. The resulting signals are then evaluated using PESQ, CSIG, COVL, CBAK, and STOI [hu2007evaluation], as summarized in Table X. We observe that PESQ exhibits only slight degradation with increasing phase shift, while CSIG, COVL, and STOI remain virtually unchanged across all shift levels. The only metric that consistently declines is CBAK, which is known to be sensitive to sample-level misalignment and likely interprets the resulting waveform discrepancies caused by a temporal shift as background noise artifacts.
@@
V-F3 Phase Reconstruction with and without PSIT
@@
For assessing the impact of PSIT on phase reconstruction quality, we conduct a controlled experiment where the model is trained to predict phase from clean ground-truth magnitude. This setup isolates the effect of phase prediction by removing confounding factors related to magnitude estimation. We compare the standard phase loss [lu2023explicit] (w/o PSIT) against four PSIT variants with increasing search grid sizes, each trained 15 times with different random seeds to ensure statistical reliability.
@@
V-F4 Training Dynamics in Multi-distortion Setting
@@
V-F5 Summary and Implications
In summary, our analyses confirm that PSIT effectively fulfills its goal of making phase supervision more tolerant to perceptually irrelevant shifts, both conceptually and empirically. As a training principle, this relaxation improves phase reconstruction quality and also facilitates more efficient and stable optimization of the entire model, including the magnitude stream. This flexibility, however, introduces minor considerations: wider search grids can increase variance in alignment-sensitive metrics such as CBAK, and a small SRMR drop is observable in the real-data dereverberation task. These effects reflect an inherent trade-off—relaxing phase supervision improves overall perceptual quality but can produce temporally shifted output waveform to which certain metrics are sensitive. Therefore, in practical use, the search grid size should be chosen based on the task goal and application, as it directly controls the amount of output shift that PSIT permits. Nonetheless, primary strength of PSIT is its practicality. It operates as a rule-based strategy that introduces only negligible computational overhead during training. It therefore offers a simple and lightweight addition to existing frameworks, yielding measurable gains in performance and efficiency without requiring architectural modifications.
@@ VII Conclusion
This paper proposes EDNet, a versatile speech enhancement framework designed to address diverse distortion types through two key components: a gated integration of masking and mapping operations, and a phase-shift-invariant training (PSIT) objective. The GM module enables the adaptive fusion of complementary enhancement strategies, while PSIT introduces robustness to temporal misalignment in phase supervision. Together, these components form a flexible and generalizable architecture for speech enhancement. We validate the proposed design through comprehensive evaluations on single- and multi-distortion tasks, where EDNet achieved state-of-the-art or comparable performance without task-specific architectural adaptation. Ablation studies and gating behavior analyses confirm that the GM module not only improves performance but also allows the model to learn task- and input-specific control over its enhancement strategy. In addition, PSIT improves phase reconstruction quality, facilitates learning dynamics, and improve overall model performance by softening the effects of strict alignment constraints without an overhead during training. These results highlight that integrating structural adaptability with shift-tolerant supervision offers a promising path toward building versatile speech enhancement systems suited for real-world deployment
VIII Acknowledgement
This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) through a grant funded by the Korean government (MSIT; RS-2025-02215122, 50%) and by the IITP–ITRC program funded by the Korean government (MSIT; IITP-2026-RS-2023-00259991, 50%).